OpenStudy (anonymous):

Medal! Casey hit the bell in the school clock tower. Her pressure reader held nearby measured the sound intensity or loudness at 40 lb/in^2 after 4 s had elapsed and at a 4.7 lb/in^2 after 7s had elapsed. She remembers from her science class that sounds decays exponentially. So the given points are (4,40) and (7,4.7) -Find an exponential equation that models these data. -How loud was the bell when it was struck at 0seconds?

OpenStudy (anonymous):

man you really have a random sample of questions this evening, huh?

OpenStudy (anonymous):

did you save them up all semester?

OpenStudy (anonymous):

hahaha kinda.. we have a lot of different topics for hw

OpenStudy (anonymous):

ok are you supposed to use something that looks like \[A=P_0e^{rt}\] or can it just be \(A=a\times b^{\frac{t}{h}}\)?

OpenStudy (anonymous):

lets go with the second one

OpenStudy (anonymous):

I thought is was just the standard exponential equation y=ab^x

OpenStudy (anonymous):

ok good, the second one

OpenStudy (anonymous):

you have two points \((4,40) \) and \( (7,4.7)\)

OpenStudy (anonymous):

first we can compute \(\frac{4.7}{40}\) as our \(b\)

OpenStudy (anonymous):

or we can do it the way we did the previous one, which is kind of tedious, but probably what they want for this problem do you have to hand this in?

OpenStudy (anonymous):

yea this is hw that we have to turn in

OpenStudy (anonymous):

ok then we do it the previous way

OpenStudy (anonymous):

but quicker

OpenStudy (anonymous):

\[7-4=3\] making \(b=\sqrt[3]{\frac{7.4}{40}}\) whatever that is

OpenStudy (anonymous):

typo there sorry \[b=\sqrt[3]{\frac{4.7}{40}}\]

OpenStudy (anonymous):

i get \(b=.49\) rounded

OpenStudy (anonymous):

then you can use \[A=40\times 4.9^{x-4}\] as your model damn this is weird

OpenStudy (anonymous):

hahahaha XD

OpenStudy (anonymous):

ok another typo, i must be getting tired \[A=40\times .49^{x-4}\]

OpenStudy (anonymous):

this is a really really dumb way to do this

OpenStudy (anonymous):

so basically you reordered the y=ab^x equation to get b=to the square root and i agree it really is

OpenStudy (anonymous):

no matter, that is what they want

OpenStudy (anonymous):

honestly you explained way more than my teacher ever did in a whole semester!

OpenStudy (anonymous):

careful, it is not the square root since \(7-4=3\) it is the cubed root

OpenStudy (anonymous):

the \(a\) in this case is what you get if \(x=4\) namely \(40\) and the \(b\) is the cubed root of the ratio \(\frac{4.7}{40}\)

OpenStudy (anonymous):

so we have \[A=40\times .49^{x-4}\] and the initial value, when \(x=0\) is \[A(0)=40\times .49^{-4}\]

OpenStudy (anonymous):

there is a much snappier way to do this much snappier

OpenStudy (anonymous):

but i don't want to confuse you, so lets leave it at that

OpenStudy (anonymous):

ok.. :) thanks again so much!! so to find how loud the bell was you have to keep it in that form?

OpenStudy (anonymous):

only an idiot starts counting at \(x=4\) a normal person starts at \(x=0\) and if you want to know what have 4 seconds early you can compute using \(-4\)

OpenStudy (anonymous):

but nvm you have the formula\[40\times .49^{x-4}\] which is what you are asked for

OpenStudy (anonymous):

ok :) yea it is a little confusing in terms of that question that they ask from you lol

OpenStudy (anonymous):

i can show you the other way to do this if you like but i don't want to make it harder for you

OpenStudy (anonymous):

sure! If you have time of course.. again thanks so much you have been really very helpful!

OpenStudy (anonymous):

btw the answer to the last question is \[40\times .49^{-4}\]

OpenStudy (anonymous):

i would do it like this start counting at \(x=0\) not \(x=4\) then since \(7-4=3\) the model is \[40\times \left(\frac{4.7}{40}\right)^{\frac{x}{3}}\]

OpenStudy (anonymous):

if i want to know what it was 4 second earlier, compute \[40\times \left(\frac{4.7}{40}\right)^{\frac{-4}{3}}\]

OpenStudy (anonymous):

oh ok makes sense :)

OpenStudy (anonymous):

Oh i see.. so basically this method is like the shortcut

OpenStudy (anonymous):

almost in any case i rounded the cube root, so it is not exact first answer is more exact

OpenStudy (anonymous):

kinda yeah i used only the numbers given \[7-4=3\] which is why my exponent was \(\frac{x}{3}\) and the ratio \(\frac{4.7}{40}\) is what i used for \(b\) rather than \(\sqrt[3]{\frac{4.7}{40}}\)

OpenStudy (anonymous):

but the other way works as well, and if you have to hand it in, use that way

OpenStudy (anonymous):

I see :) thanks again!! This really helped a lot hope you enjoy the rest of your day ^.^

OpenStudy (anonymous):

you too i hope you are done with the random questions (for your sake) good luck

OpenStudy (anonymous):

haha yes all done :) now I just have to prep for a test next week! :)